OpenStudy (anonymous):

***HELP*** A car accelerates uniformly from +10.0 m/s to +40.0 m/s over a distance of 125 m. How long did it take to go that distance? Show all your work, including the equation used, given and unknown quantities, and any algebra required. Make sure your answer has the correct number of significant figures.

OpenStudy (ivancsc1996):

\[v _{f}^{2}=v _{i}^{2}+2ax=1600\frac{ m ^{2} }{ s ^{2} }=100\frac{ m ^{2} }{ s ^{2} }+2a125m \rightarrow a=\frac{ 1600\frac{ m ^{2} }{ s ^{2} }-100\frac{ m ^{2} }{ s ^{2} } }{ 2 \times 125m }\]Since \[a=\frac{ v _{f}-v _{i} }{ t }\]Then\[t=\frac{ (40\frac{ m }{ s }-10\frac{ m }{ s })125m2 }{ 1600\frac{ m ^{^{2}} }{ s ^{2} }-100\frac{ m ^{2} }{ s ^{2} } }=5s\]